during a game, the pittsburgh steelers converted 4 of 12 fourth downs into first downs. with only minutes left in the game, it is fourth down again and they must make a first down to continue their drive. what is the probability that they will make the first down?
Answer by stanbon(75887)

Prob(success)=4/12 = 1/3
Cheers,
Stan H.